A liquid fertilizer requires dilution before use. How much liquid fertilizer must be added to 600mL of water to make a 25% fertilizer solution? Can you show the work please?
Answer by ptaylor(2198)

Let x=amount of liquid fertilizer that needs to be added
Now we know that the amount of pure fertilizer that exists before the mixture takes place has to equal the amount of pure fertilizer that exists after the mixture takes place, sooooo:
x=0.25(600+x)
x=150+0.25x subtract 0.25x from each side
x-0.25x=150
0.75x=150
x=200 ml----amount of liquid fertilizer needed
CK
200ml=0.25*800ml
200ml=200ml
Hope this helps---ptaylor